Determine ho and Side flow rate and the power loss using the following

given parameters: Temperature = 62 C, N = 40 rev/s, W = 2 kN (bearing

load), r = 20 mm, c = 0.04mm, and l = 3.2 cm


1
Expert's answer
2021-08-20T01:36:54-0400

velocity

"V = \frac{2\pi N}{60} =\frac{2\pi 40}{60} = 4.19 \frac{m}{sec}"


flow rate

"Q = \frac{\pi}{4}D^2 V = \frac{\pi}{4}\times 0.04^2 \times4.19 =0.0053 \frac{m^3}{sec}"


Darcy’s equation to calculate the loss 

"h_L=f\times \frac{L}{D}\times \frac{v^2}{2g}
\,."

"h_L=0.00004\times \frac{0.032}{0.04}\times \frac{4.19^2}{2\times9.81} = 2.86\times10^-5 m =0.0286mm
\,."